OpenStudy (anonymous):

Marisol plays a game using the spinner shown. During the game she will spin a total of 54 times. Which is the best prediction for the number of times she will get either "forward 1" or "forward 2" on the spinner?

OpenStudy (anonymous):

@Dbzfan836

OpenStudy (anonymous):

@johnweldon1993 @jim_thompson5910

OpenStudy (johnweldon1993):

"Spinner shown" Is there a diagram?

OpenStudy (anonymous):

@johnweldon1993

OpenStudy (anonymous):

@ciaran.awesome

OpenStudy (anonymous):

@cwrw238

jimthompson5910 (jim_thompson5910):

how many spaces are there to land on?

OpenStudy (anonymous):

6

jimthompson5910 (jim_thompson5910):

how many spaces are "forward 1" or "forward 2"

OpenStudy (anonymous):

what do you mean?

jimthompson5910 (jim_thompson5910):

do you see the spaces labeled "forward 1" or "forward 2"

OpenStudy (anonymous):

yes i do

jimthompson5910 (jim_thompson5910):

how many of those are there?

OpenStudy (anonymous):

1 of each and 2 together

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

2 spaces that we want, out of 6 total

jimthompson5910 (jim_thompson5910):

so 2/6 = 1/3 is the probability of landing on either one

jimthompson5910 (jim_thompson5910):

multiply this probability by the number of spins to get your answer

OpenStudy (anonymous):

so what is the problem? multiply 2/6*1/3 ?

jimthompson5910 (jim_thompson5910):

no

jimthompson5910 (jim_thompson5910):

you'll multiply 1/3 by the number of spins

OpenStudy (anonymous):

so is the answer 2?

jimthompson5910 (jim_thompson5910):

how many spins are there total?

OpenStudy (anonymous):

its 54 times its spins, so the answer is 18 right?

jimthompson5910 (jim_thompson5910):

54*(1/3) = 18, yep

jimthompson5910 (jim_thompson5910):

you expect to land on those spaces about 18 times